C and Y candy company mixes candy that costs $6.00 per kg with candy that costs $4.50 per kg. How many kg's of each are needed to make a 3kg box that costs $15.00 ?
Answer by checkley77(12569)

6x+4(3-x)=15
6x+12-4x=15
2x=15-12
2x=3
x=3/2
x=1.5 kg of the $6 candy is used.
3=1.5=1.5 kg of the $4 candy is used.
Proof:
6*1.5+4*1.5=15
9+6=15
15=15